1) A Gaussian random variable has a mean value of 3 and a standard deviation of 2. Find the probability that the value of the random variable exceeds 9. Repeat for the probability that it is less than -5.

ANSWER WITH COMPLETE STEPS THANKS
Answer #1

1)
(a)

P(x > 9) = 1 - P(x < 9)

= 1 - P[(x - ) / < (9 - 3) / 2]

= 1 - P(z < 3)

= 0.0013

Probability = 0.0013

(b)

P(x < -5) = P[(x - ) / < (-5 - 3) / 2]

= P(z < -4)

= 0

Probability = 0